OpenStudy (anonymous):

MEDALS GIVEN Which ordered pair is a solution to the inequality y > x + 5? (-3, 2) (0, -3) (3, 5) (2, 8)

OpenStudy (anonymous):

What is the slope between the two points (8, 15) and (3, 5)?

OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2? y = 2x + 1 y = 2x + 3 y = 2x - 8 y = 2x - 3

OpenStudy (anonymous):

you could just substitute the values to see which one satisfies right?

OpenStudy (anonymous):

Im just looking for the answer i really need it fast.

OpenStudy (anonymous):

substituting the values on the coodinates will take you no time.

OpenStudy (anonymous):

that's a very simple linear equation.

OpenStudy (anonymous):

on the first one

OpenStudy (anonymous):

I dont know how to do that.

OpenStudy (anonymous):

in the coordinate, the first number is the x and the 2nd number is the y (x,y)

OpenStudy (anonymous):

so, just try to look which one satisfies =))

OpenStudy (anonymous):

in finding the slope, do you know the 2-point form formula?

OpenStudy (anonymous):

just substitute the first point's x to the x sub and the 2nd point's x to the x sub 2. same goes with y.

OpenStudy (anonymous):

on the 3rd, just substitute the coordinate in all the equations to see what is correct